-- SimoneEsposito - 16 Sep 2013

Daytime Activity

The daytime activity started at 9:00 after a call with Itlay to discuss the today plan.

Quick check of pupil wobble at BIN1 (pupil TN 184442)

We do a quick check of the pupil wobble with the BIN1. We plan to do a more detailed test but we want first to measure BIN2 to be ready to test it on sky tonight if possible.
Inst rot Cy Cx  
300 82 0  
210 30 32  
120 -8 -17  
30 50 -50  
-40 100 0  
The camera lens is not able to recenter the pupil when instrument rotator is around -30 degree. The stroke of Y axis is not sufficient. We implement a digital shift of one pixel with the script movePupils. We start from the pupil TN 184442. We generate a new set of pupil with TN 20130916_165012.

We close the loop with the new pupil to check we desaturated the Y axis.

Instrument rotator at -30. The camera lens Y axis is now Y=67, X=-26.

Instrument rotator at -60. The camera lens Y axis is now Y=64, X=-3.

Instrument rotator at 30. The camera lens Y axis is now Y=20, X=-52

The check is positive and we accept the ne pupils with TN 20130916_165012

BIN2 (10:07)

We adjust the rotation for the BIN2.

The rotation offset of the puil rerotator is 33.5 degree as found from previous files. We add to the new BIN1 offset 33.5 degree to get the BIN2 rerotator offset.

NOTE: because of the new offset between the instrument rotator and pupil derotator the pupil wobble with BIN2 will be different from BIN1. So it will need a second pupil wobble check.

We create a pupil with a single subaperture, the central one, removed. See the estimated geometry in the image below. The new BIN2 pupil has a TN 173526

BIN2 pupil 173526 (1 central subaperture masked)

IM measurements with KL_V16

We take a measurements of a 153 modes IM with the new pupils (173526). measurement is done with 48 push pull average.

BIN2 main parameters => 3lambda/D 1; 53modes

IM 174226

REC 174622

BIN2 wobble test

We test the pupil whobble at BIN2. The results is acceptable. results are reported in the table below.

<table border="1" cellpadding="0" cellspacing="1">

intrument rot | cy | cx || 300 | -67| 6| 210| -56| 38| 170| -66 120| -30 | -12 30| 31| 19 -30| 8| 63 -60| -16| 79 -90| -34| 82

Test of BIN2 rec in closed loop without/with disturbance hig flux

we optimize the gain with the autogain. No force peaks seen.
Test in closed loop with disturbance at 0.8 arcsec. 1000fps mag=7
TN 175538
we decide to test changin the modal gain groups. The new groups are 2, 80. They were 2, 30.

Test in closed loop with disturbance 0.8

TN 180534 ->>> 181446 IRTC SR H band range 55-59

test closed loop without disturbance

TN 184142 ->>> 184607 IRTC SR H band range 74-80

BIN2 test of closed loop disturbance mag 10.7 0.8 arcsec

optimized gains 0.7, 0.5, 0.4

TN 184442 ->>> 185925 IRTC SR H band 47-54

We see a vibration line at 60 Hz.

We teake some TN to see if the 60Hz vibration has a time periodicity.

TN 190805 ->>> 191315 IRTC SR H band 45-55

Digital shift of the pupil

we discover that the camera lens do not cover the full range of rotation we try to do a digital shift of one pixel. The new pupil has a TN 193614. With this adjustment we

BIN2 test of closed loop disturbance mag 11.7 0.8 arcsec

we test the BIN2 at a dimmer magnitude. The closed loop diverges. One edge of the mirror is lost. It seems due to a low order mode. The patch of lost edge is the one where we have 2/3 disabled actuators. We noted that we miss a few subaperture on that edge. We modified the pupil mask to include 2 subapertures more. This is done to see if we are able to control that edge better.

a) adding two subapertures did not chenge the beaviour

b) removing 1 subaperture on the opposite side of the edge.

We discover that the considered edge was vignetted by the beam splitter cube. So we start again.

New pupil BIN2 after BS adjust

We acquire a new pupil set changin the figure of merit of the hottization procedure. We select the subapertures now using the intensity threshold. We select all the subaperture with 0.28 of the average intensity.

New pupil TN ->>> 220247

New BIN2 IM and REC with new pupils

153 modes 3lmbda/D

IM 220905

REC 221257

Test at high flux is positive no edge divergence is found.

The lamp is set to mag 11.autogain is run, OK, Closed loop test is repeated. Photons/subaps 120. Test is OK.

The lamp is set to mag 11.7, autogain is run, OK. Test is repeated. Photons/subaps is 60. Test is OK.

We acquire PSF for this magnitudes. 625fps, 3lambda/D BIN2

TN 222507 ->>> 222537

TN 222901 ->>> 222931

SR values comprised between 35%-40%

BIN2 test at high flux

REC 221257

The lamp is set to mag 7.9. Autogain is run, OK, Closed loop test is repeated. Photons/subaps 1300. disturbance 0.8

6 TN hig flux 224003 ->>> 224350

SRs in range 56-58.

Test of Agapito filters

We find a problem in loading the reconstructor matrix with the IIR filter. The secondary ripped. The test is stopped for the telescope hand over.

Figures

WFS subaps mask geometry
  • metry: 2013-09-16_10.33.48.jpg
Topic attachments
I Attachment Action Size Date Who Comment
2013-09-16_10.33.48.jpgjpg 2013-09-16_10.33.48.jpg manage 94 K 16 Sep 2013 - 18:41 UnknownUser WFS subaps mask geometry
Topic revision: r5 - 17 Oct 2014, AlfioPuglisi
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback